OpenStudy (anonymous):

Does anyone know how to find the derivative of f(x)=(sqrt(1/x^7)) the square root over the whole thing is confusing me

hartnn (hartnn):

do you know the derivative of x^n ?

OpenStudy (anonymous):

no I dont think so

hartnn (hartnn):

\(\large \dfrac{d}{dx}x^n=n \times x^{n-1}\)

OpenStudy (anonymous):

well I know that if the square root wasnt over the whole problem, then wouldn't the answer be -7/x^6? I just am confused on the square root being over the whole thing

hartnn (hartnn):

how did u get -7/x^6 ?

hartnn (hartnn):

\(\huge \sqrt A = A^{\frac{1}{2}}\)

OpenStudy (anonymous):

\[\sqrt{(1)/(x^7)}\]

OpenStudy (anonymous):

so if the square root wasnt there wouldnt the answer be -7x^-6?

hartnn (hartnn):

i mean you have used the formula that i gave you to find derivative of x^7 , right ? but since we have square root, it'll be \(\large x^{7/2}\)

hartnn (hartnn):

in the denominator

hartnn (hartnn):

so, basically you are differentiating, 1/x^(7/2) or \(\large x^{-7/2}\)

OpenStudy (anonymous):

thats the answer? I'm sorry I'm just confused lol

hartnn (hartnn):

thats what you will be differentiating. what will u take 'n' as ?

OpenStudy (anonymous):

so it'd be -7/2x^-4.5?

OpenStudy (anonymous):

\[-7/2x^9/2\]

hartnn (hartnn):

yes! but keep -9/2 as it is not 4.5

OpenStudy (anonymous):

Thank you so much! :)

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

since you are new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) if you have any doubts browsing this site, you can ask me. Have fun learning with Open Study! :)

OpenStudy (anonymous):

Okay thanks again! :)